OpenStudy (anonymous):

If m = 7x + 7, m = 4y, and m = 112, find the values of x and y

OpenStudy (anonymous):

112=7x+7 -7 105=7x /7 15=x --- 112=4y /4 28=y

OpenStudy (anonymous):

x=15.. and y=28..

OpenStudy (anonymous):

@ashleyyymyerss does that help??

OpenStudy (thomas5267):

Split it into two steps. First, find \(x\) using \(112=7x+7\). The use \(x\) and \(4y=7x+7\) to find \(y\).

OpenStudy (anonymous):

so if m=112.. for the first one its 112=7(15)..second one would be 112=4(28).. you should get 112 for both..